Need help in solving this it is not a test but for our homework. Every way I have worked it I get a different answer. 0.4p+0.2=4.2p-7.8-0.6p
Answer by stanbon(75887)

0.4p+0.2 = 4.2p-7.8-0.6p
0.4p+0.2 = -0.2p-7.8
0.6p = -8
p = -8/(3/5)
p = -40/3 = -(13 1/3)

===============
Cheers,
Stan H.